# No Representation, No Trust: Connecting Representation, Collapse, and Trust Issues in PPO

### v1

@@ 4 Intervening to regularize representations and non-stationarity
Asadi et al. (2023) argue that as the targets of the value function change with the changing policy rollouts, the old moments accumulated by Adam become harmful to fit the new targets and find that resetting the moments of Adam helps performance in DQN-like algorithms. As the PPO objective creates a dependency on the previous policy, and more generally, in the policy gradient, the advantages change with the policy, the same argument about Adam moments can be made for PPO. Therefore, we experiment with resetting Adam’s moments after each batch collection (to avoid tuning its frequency) for both the actor and the critic; the moments are thus only accumulated over the epochs on the same batch. We observe that this intervention reduces the feature norm and increases the feature rank on ALE, which also reduces the excess probability ratio and, in some cases, improves performance; however, it is not as effective as the other interventions in preventing collapse and, like sharing the actor-critic trunk results in poor performance on ALE/Gravitar. Nevertheless, further tuning the frequency of the reset or coupling it with representation dynamics may yield different conclusions.
@@ Conclusion
In this work, we have provided evidence that the representation deterioration under non-stationarity observed by previous work in value-based methods generalizes to PPO agents in ALE and MuJoCo with their common model architectures and can lead to performance collapse. We have shown that this is particularly concerning for the heuristic trust region set by PPO-Clip, which fails to prevent collapse as it becomes less effective when the agent’s representation becomes poor. Finally, we presented Proximal Feature Optimization (PFO), a simple novel auxiliary loss based on regularizing the features’ evolution, which mitigates representation degradation and along with other interventions shows that controlling the representation degradation improves performance.
@@
Although we study pixel-based discrete action environments and continuous control environments, we only study the common architectures of each environment, which do not include normalization layers and are limited to relatively small architectures with no memory (i.e., exclude Transformers or RNNs). Therefore, the generalization of the results of this work to other settings is unknown. Still, this work should raise awareness about the representation collapse phenomenon observed in PPO and encourage future work to monitor representations when training PPO agents, as failure to train indefinitely and sudden collapse may be due to representations. We have presented here methods that mitigate representation degradation, but none of them seem to be a fundamental solution and a deeper understanding of the reasons driving representation deterioration under non-stationarity is still needed. We plan to address these issues in future work.
@@ B.1 Code and run histories
Our code is available at https://github.com/CLAIRE-Labo/no-representation-no-trust. It includes the development environment distributed as a Docker image for GPU-accelerated machines and a Conda environment for MPS-accelerated machines, the training code, scripts to run all the experiments, and the notebook that generated the plots. The codebase uses TorchRL (Bou et al., 2024) and provides a comprehensive toolbox to study representation dynamics in policy optimization. We also provide modified scripts of CleanRL (Huang et al., 2022b) to replicate the collapse observed in this work and ensure it is not a bug from our novel codebase.
@@
Runs are fully reproducible on the same acceleration device on which they were run.
@@ B.2 Additional details on our experimental setup
Num envs in parallel plasticity 1
@@
Total agent steps plasticity 36,000 (at least one full episode)
@@
Number of epochs plasticity fit 1
@@
Plasticity every 2.5% (41 times in total)
Table 1: Hyperparameters for ALE.
@@
Num envs in parallel plasticity 4
@@
Total environment steps plasticity 4,000 (at least 4 full episodes)
@@
Number of epochs plasticity fit 4
@@
Plasticity every 2.5% (41 times in total)
Table 2: Hyperparameter for MuJoCo.
To measure the plasticity loss of a checkpoint, we use the same optimization hyperparameters used to train the checkpoint, i.e. the same batch size and learning rate. The dataset sizes and fitting budgets for plasticity are listed in Tables 1 and 2.
We provide a high-level pseudocode for PPO in Algorithm 1 and list all hyperparameters considered in Tables 1 and 2.
Algorithm 1 High-level Pseudocode for PPO
1:$N$: number of environments in parallel.
2:$B_{\text{env}}$: agent steps per environment to collect in a batch.
3:$K$: number of optimization epochs per batch.
4:
5:$L_{\pi_{\text{old}}}^{\text{CLIP}}({\bm{\theta}})$: PPO-Clip objective.
8:
9:while collected environment steps $\leq$ total environment steps do
11:     for $\text{actor}=1$ to $N$ do
14:     end for
15:     Minimize overall policy and value loss $(-L_{\pi_{\text{old}}}^{\text{CLIP}}({\bm{\theta}})-c_{H}H({\bm{\theta}})+c_{VF}L^{VF}({\bf w}))$ with autograd on the on the collected batch over $K$ epochs with minibatch size $M\leq B$.
16:     $\pi_{\text{old}}\leftarrow\pi_{\bm{\theta}}$
17:end while
Proximal Feature Regularization
With a coefficient $c_{PFO}$, the PFO loss is added to the overall loss $(-L_{\pi_{\text{old}}}^{\text{CLIP}}({\bm{\theta}})+c_{PFO}L_{\pi_{\text{old}}}^{PFO}({\bm{\theta}})-c_{H}H({\bm{\theta}})+c_{VF}L^{VF}({\bf w}))$ optimized with autograd over multiple mini-batch epochs.
@@ B.3 Additional details on metrics used in the figures
We group the different epoch configurations of an intervention on the same environment, giving 15 runs per boxplot (three epochs with five seeds each).
B.4 Hardware and runtime
The experiments in this project took a total of 10,368 hours between NVIDIA V100 and A100 GPUs (ALE) and CPUs (MuJoCo). A run on ALE takes around 12 hours on an A100 and 24 hours on a V100. A run on MuJoCo takes around 5 hours on CPUs.
@@ Appendix D Main paper figures on all environments
Figure 7: Figure 1 on ALE. QBert is the only game where rank decline and collapse are not observed, apart from an outlier run that collapsed at initialization.
@@
Figure 18: Figure 6 on ALE. The tails of the plasticity loss on Phoenix with interventions can be higher than without interventions on the runs where the models collapse too early without interventions, leading to the plasticity loss of the non-collapsed models with interventions eventually becoming higher. This can be observed from the training curves with interventions. Nevertheless, their medians are lower.

### v3

@@ 4 Intervening to regularize representations and non-stationarity
Asadi et al. (2023) argue that as the targets of the value function change with the changing policy rollouts, the old moments accumulated by Adam become harmful to fit the new targets and find that resetting the moments of Adam helps performance in DQN-like algorithms. As the PPO objective creates a dependency on the previous policy, and more generally, in the policy gradient, the advantages change with the policy, the same argument about Adam moments can be made for PPO. Furthermore, Dohare et al. (2023b); Lyle et al. (2023) advocate for decaying the second moment of Adam faster than its default decay of 0.999 when training under non-stationarity and set it to match the decay of the first moment. Therefore, we experiment with both resetting Adam’s moments after each batch collection (to avoid tuning its frequency) and setting the second moment to decay at the (smaller) default decay of the first moment for both the actor and the critic; the moments are thus only accumulated over the epochs on the same batch in the former and over shorter batch sequences in the latter. We observe in Figure 6 and Appendix D that these interventions reduce the feature norm and increase the feature rank on ALE, which also reduces the excess probability ratio and, in some cases, improves performance; however, they are not sufficient to prevent collapse and, like sharing the actor-critic trunk, result in poor performance on ALE/Gravitar.
5 Related Work
Our work is complementary to various other works studying the plasticity and representation dynamics of neural networks trained under non-stationarity. Kumar et al. (2023) provide a comprehensive comparison and categorization of methods used to mitigate plasticity loss in continual supervised learning tasks and their effects on representations. Our work provides insights into the transferability of some of these solutions to RL and tools to evaluate their impact on trust region methods. Sokar et al. (2023) provide an alternative characterization of plasticity loss in RL using dormant neurons and observe an increase in dormant neurons for non-stationary objectives. Abbas et al. (2023) study representation metrics such as feature norms and observe a decrease of the norm due to dying neurons. Like in the work of Lyle et al. (2022), both studies only include value-based methods. In this work, we study dead units and capacity loss as Lyle et al. (2022) and provide corroboration of the dying units phenomenon in policy optimization methods and, taking the dying neurons out of the equation, find that the norm of preactivations actually blows up.
Other feature regularizations similar to PFO have been studied in value-based offline RL. Kumar et al. (2022) propose DR3, which counteracts an implicit regularization in TD learning by minimizing the dot product between the features of the estimated and target states. Ma et al. (2023) propose Representation Distinction (RD) which tries to avoid unwanted generalization by minimizing the dot product between the features of state-action pairs sampled from the learned policy and those sampled from the dataset or an OOD policy. Both are related to PFO as the methods directly tackle an undesired feature learning dynamic, but there is no motivation for DR3 or RD in online RL, and PFO is conceptually different. The implicit regularization that DR3 counteracts is not present in on-policy RL as shown by Kumar et al. (2022) in the SARSA experiment, and PFO differs from DR3 as it extends a trust region rather than counteracts an implicit bias. Similarly, the overestimation studied by Ma et al. (2023) in the vicious backup-generalization cycle is broken by on-policy data, and RD regularizes state features between the learned policy and the dataset policy, not consecutive policies.
@@ Conclusion
In this work, we provide evidence that the representation deterioration under non-stationarity observed by previous work in value-based methods generalizes to PPO agents in ALE and MuJoCo with their common model architectures and is connected to performance collapse. This brings a novel perspective to previous works that showed that PPO agents lose plasticity throughout training. Weshow that this is particularly concerning for the heuristic trust region set by PPO-Clip, which fails to prevent collapse as it becomes less effective when the agent’s representation becomes poor. Finally, we present Proximal Feature Optimization (PFO), a simple novel auxiliary loss based on regularizing the evolution of features that mitigates representation degradation and, along with other interventions, shows that controlling representation mitigates performance collapse.
@@
In this work, we study the common architecture and optimizer of PPO agents in ALE and MuJoCo consisting of relatively small models without normalization layers, weight decay, or memory (e.g., not using Transformers and RNNs). Despite our best attempts, as with any other empirical machine learning work, the generalization of our results to other settings is not fully known. Still, this work should raise awareness about the representation collapse phenomenon observed in PPO and encourage future work to monitor representations when training PPO agents, as it can help diagnose performance collapse. We have focused on simple interventions that regularize non-stationarity and representations to highlight the effects of non-stationarity and the connection between representation, trust region, and collapse, but exploring interventions on plasticity is also valuable, as these may also influence the same dynamics. We believe further studies to analyze this problem, both empirically and particularly theoretically, to understand the reasons driving representation deterioration to be valuable. We hope that our study encourages work in this direction.
Acknowledgments and Disclosure of Funding
We extend our gratitude to the reviewers for their valuable insights, which significantly enhanced the clarity and rigor of this work. We are particularly grateful to the area chair for their guidance which shaped the final version of this paper. We thank the SCITAS team at EPFL for the access to the beta testing phase of their new cluster. We are also grateful to Vincent Moens for his support with the TorchRL library. Finally, we thank nimble.ai for their generous gift to the CLAIRE Lab, which supported D.P.’s Master’s project.
@@ B.1 Code and run histories
Our codebase is publicly available at https://github.com/CLAIRE-Labo/no-representation-no-trust. It includes the development environment distributed as a Docker image for GPU-accelerated machines and a Conda environment for MPS-accelerated machines, the training code, scripts to run all the experiments, and the notebook that generated the plots. The codebase uses TorchRL (Bou et al., 2024) and provides a comprehensive toolbox to study representation dynamics in policy optimization. We also provide modified scripts of CleanRL (Huang et al., 2022b) to replicate the collapse observed in this work and ensure it is not a bug from our novel codebase.
@@
Runs are fully reproducible on the same acceleration device on which they were run. In particular, we have reproduced our results on three different clusters with the same NVIDIA GPU device.
@@ B.2 Additional details on our experimental setup
To measure the capacity loss of a checkpoint, we use the same optimization hyperparameters used to train the checkpoint, i.e. the same batch size and learning rate. The optimizer is reconstructed from its initial state (loading the optimizer state is also a valid design choice). The dataset sizes and fitting budgets for capacity are listed in Tables 2 and 3.
We provide a high-level pseudocode for PPO in Algorithm 1 and list all hyperparameters considered in Tables 2 and 3.
Algorithm 1 High-level Pseudocode for PPO
1:$N$: number of environments in parallel.
2:$B_{\text{env}}$: agent steps per environment to collect in a batch.
3:$K$: number of optimization epochs per batch.
4:
5:$L_{\pi_{\text{old}}}^{\text{CLIP}}({\bm{\theta}})$: PPO-Clip objective.
8:
9:while collected environment steps $\leq$ total environment steps do
11:     for $\text{actor}=1$ to $N$ do
14:     end for
15:     Minimize overall policy and value loss $(-L_{\pi_{\text{old}}}^{\text{CLIP}}({\bm{\theta}})-c_{H}H({\bm{\theta}})+c_{VF}L^{VF}({\bf w}))$ with autograd on the collected batch over $K$ epochs with minibatch size $M\leq B$.
16:     $\pi_{\text{old}}\leftarrow\pi_{\bm{\theta}}$
17:end while
Proximal Feature Regularization
With a coefficient $c_{PFO}$, the PFO loss is added to the overall loss $(-L_{\pi_{\text{old}}}^{\text{CLIP}}({\bm{\theta}})+c_{PFO}L_{\pi_{\text{old}}}^{PFO}({\bm{\theta}})-c_{H}H({\bm{\theta}})+c_{VF}L^{VF}({\bf w}))$ optimized with autograd over multiple minibatch epochs.
Table 1: Hyperparameters for the toy setting in Figure 5.
Environment
$\phi(x)$ Sampled from a Normal distribution
$\phi(y)$ $\alpha\phi(x)$
$\alpha$ 3 (overshoot), -1 (interfere)
$A(x,a_{1}),A(y,a_{1})$ 1
Policy
Network 2 output neurons representing the 2 logits + Softmax
Optimization
Clipping epsilon (PPO-Clip) 0.1
Optimizer SGD
Learning rate 1.5
Minibatch size 1
Number of epochs 10
Number of steps 20 alternating between $x$ and $y$
Table 2: Hyperparameters for ALE.
@@
Num envs in parallel capacity 1
@@
Total agent steps capacity 36,000 (at least one full episode)
@@
Number of epochs capacity fit 1
@@
Capacity every 2.5% (41 times in total)
Table 3: Hyperparameters for MuJoCo.
@@
Num envs in parallel capacity 4
@@
Total environment steps capacity 4,000 (at least 4 full episodes)
@@
Number of epochs capacity fit 4
@@
Capacity every 2.5% (41 times in total)
@@ B.3 Additional details on metrics used in the figures
Training curves
A point in the training curves in Figures 1, 2, 3, before aggregating seeds and smoothing, corresponds to an average value over the last batch collected at the time of logging for metrics available at every batch (feature rank, entropy, etc.) or the latest batch where the metric was available at the time of logging for the episodic return (as it’s only available when episodes finish, and it requires multiple batches to finish an episode). E.g., in Figure 1 on ALE, a feature rank corresponds to the average feature rank over all the states in the last batch collected at the time of logging and is logged every 0.1% of the batches (i.e. every 6,144 env steps); A return corresponds to the average return across all workers that had episodes finished in the latest batch containing finished episodes at the time of logging.
@@
We group the different epoch configurations of an intervention on the same environment, giving 15 runs per boxplot (three epochs with five seeds each). The right (resp. left) whiskers are determined by the highest (resp. lowest) observed datapoint below Q3 + 1.5 IQR (default of Matplotlib). The outliers are points outside of the whiskers.
B.4 Statistical significance
Stochasticity in our experiments arises from network initialization, environment transitions (e.g., sticky actions in ALE), agent action sampling, and minibatch sampling for optimization. A seed fully controls the sequence of randomness in a run with the same hyperparameter configuration. We repeat each configuration with five seeds using the same collection of seeds, resulting in the same initialization of the networks and environments for a given seed across configurations. This form of repeated measures allows us to compare the configurations with lower variance as they share the same initial conditions, hence requiring a lower number of seeds.
In Figures 1 and 2, we aggregate the five runs of each experiment into mean curves with min/max shaded areas. The use of min/max error bars allows us to demonstrate the full range of observed outcomes, although it may result in shaded areas that overlap more than with other types of error bars. Most of the claims we make based on those figures do not rely on non-overlapping shaded areas and are instead stronger when the max or min boundaries are consistent with the observation made (min boundary of feature norm increasing, max boundary of feature rank decreasing). Otherwise, we made comparative claims when shaded areas did not overlap (feature rank decreasing faster with more epochs and more non-stationarity).
Figure 3 displays individual seeds to zoom on single-run dynamics around collapse. It is used for an illustrative purpose to provide intuition and does not depend on the number of runs or statistical aggregation of results. The main claim made with the intuition (breakdown of the trust region) is backed by Figure 4, which includes 300 points per plot per environment, subsampled from 15 training curves per environment.
To evaluate the effects of the interventions in Figure 6, we show boxplots to give a complete idea of the distribution of the data which is formed by grouping the different configurations in the same environment. Each boxplot contains 15 runs. We make claims such as preventing collapse using the tails and medians and claims about lower excess ratio and higher rank using the interquartile range. Without a clear intuition about the distribution of combined configurations per environment, we consider this approach appropriate for comparing interventions.
In summary, we believe our experimental design provides a balanced tradeoff between statistical significance and richness of claims. The computational cost of running more seeds may not yield proportionately valuable insights.
B.5 Hardware and runtime
The experiments in this project took a total of ~11,300 GPU hours on NVIDIA V100 and A100 GPUs (ALE) and ~25,500 CPU hours (MuJoCo). A run on ALE takes around 10 hours on an A100 and 16 hours on a V100. A run on MuJoCo takes around 5 hours on 6 CPUs.
@@ Appendix D Main paper figures on all environments
Figure 7: Figure 1 on ALE. QBert is the only game where rank decline and collapse are not observed, apart from an outlier run that collapsed at initialization. The performance of the policy should be taken into consideration when comparing the capacity loss of the critic. E.g., for Phoenix, the capacity loss of the critic associated with the policy that doesn’t collapse ends up higher than that of the policies that do collapse.
@@
Figure 18: Figure 6 on ALE. The tails of the capacity loss on Phoenix with interventions can be higher than without interventions on the runs where the models collapse too early without interventions, leading to the capacity loss of the non-collapsed models with interventions eventually becoming higher. This can be observed from the training curves with interventions. Nevertheless, their medians are lower.
@@
Figure 26: Figure 1 on MuJoCo HalfCheetah with the tanh activation.
Figure 27: Figure 1 on MuJoCo Ant with the tanh activation.
@@
Figure 30: Figure 1 on MuJoCo HalfCheetah with the ReLU activation.
Figure 31: Figure 1 on MuJoCo Ant with the ReLU activation.
Figure 32: PPO on ALE/Phoenix-v5 collapses with its standard tuned hyperprameters from (Schulman et al., 2017) (4 epochs) when training for 200M steps. Regularizing with PFO mitigates the collapse (applied on the last pre-activation).
Figure 33: Three seeds of PPO on ALE/NameThisGame-v5 showing that it also collapses with its standard tuned hyperprameters from (Schulman et al., 2017) (4 epochs) when trained for long enough.
Figure 34: Varying $\varepsilon$ cannot be used as a reliable tool to study collapse as monotonic changes in $\varepsilon$ yield non-monotonic collapse speeds, unlike when varying the number of epochs.
Figure 35: PPO with a shared actor-critic on ALE/Phoenix-v5 with a sparse reward (random masking with probability 0.9) with the standard hyperparameters (4 epochs). While with dense rewards, sharing the trunk was beneficial in ALE/Phoenix (Figure 21 Appendix), with the sparse reward, the opposite is true: sharing the trunk is detrimental.
@@ Results
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Using the numbered contributions and claims in the introduction, we provide evidence for claims 1, 2, 3, 4, and 5 in Sections 3.1, 3.2, 4, 4, and Appendix B.1, respectively.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are discussed in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Proofs relative to Section 3.2.1 are presented in Appendix C.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We list all the experimental details in Appendix B and provide a fully reproducible codebase.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide a fully reproducible codebase with instructions in Appendix B.1.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe our experimental setup in Section 3 and give details in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We discuss statistical significance in Appendix B.4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We discuss computing resources in Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper addresses a fundamental algorithmic issue with no specific application. The impact of any application should be discussed specifically when introducing the application.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The codebase implements already existing algorithms and training curves do not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the assets used have been cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We introduce a codebase and run histories documented in Appendix B.1.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing or research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing or research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.